# From Voice to Safety: Language AI Powered Pilot-ATC Communication Understanding for Airport Surface Movement Collision Risk Assessment

### v1

@@ Abstract
Surface movement collision risk models are critical to enhancing airport safety. These models play a vital role in identifying and mitigating potential hazards during ground operations by providing early warnings for near-miss incidents, thereby reducing the risk of accidents that could jeopardize human lives and financial assets. However, existing models, developed decades ago, have not fully integrated recent advancements in machine intelligence, where incorporating additional functionalities presents promising opportunities for improved risk assessment. This work provides a feasible solution to the existing airport surface safety monitoring capabilities (i.e., Airport Surface Surveillance Capability (ASSC)), namely language AI-based voice communication understanding for collision risk assessment. The proposed framework consists of two major parts, (a) Automatic Speech Recognition (ASR); (b) surface collision risk modeling. ASR module generates information tables by processing voice communication transcripts, which serve as references for producing potential taxi plans and calculating the surface movement collision risk. For ASR, we collect and annotate our own Named Entity Recognition (NER) dataset based on open-sourced video recordings and safety investigation reports. Additionally, we refer to FAA Order JO 7110.65W and FAA Order JO 7340.2N to get the list of heuristic rules and phase contractions of communication between the pilot and the Air Traffic Controller (ATCo) used in daily aviation operations. Then, we propose the novel ATC Rule-Enhanced NER method, which integrates the heuristic rules into the model training and inference stages, resulting into hybrid rule-based NER model. We show the effectiveness of this hybrid approach by comparing different setups with different token-level embedding models. For the risk modeling, we adopt the node-link airport layout graph from NASA FACET and model the aircraft taxi speed at each link as a log-normal distribution and derive the total taxi time distribution. Then, we propose a spatiotemporal formulation of the risk probability of two aircraft moving across potential collision nodes during ground movement. We show the effectiveness of our approach by simulating two case studies, (a) the Henada airport runway collision accident happened in January 2024; (b) the KATL taxiway collision happened in September 2024. We show that, by understanding the pilot-ATC communication transcripts and analyzing surface movement patterns, the proposed model estimate the surface movement collision probability within machine processing time, thus enables proactive measures to runway incursions, taxiway collisions, and thus improve airport safety by providing risk assessment in time. A study on validating the log-normal assumption of aircraft taxi speed distributions is also given. The link to code and data repository for this research work can be found from this link.
@@
We introduces a novel framework to integrate natural language processing with surface collision risk modeling to enhance aviation safety.
We provide a way to create and annotate the pilot-ATC communication voice for downstream NLP tasks.
@@
Surface movements is modeled with log-normal distributions and employed node-link graph structures to estimate spatiotemporal collision probabilities.
We valid the log-normal link travel speed assumptions by conducting data analysis and statistical tests on ASDE-X ground movement data.
The simulation with the Haneda runway collision and the KATL taxiway collision case studies validated the effectiveness on detecting high-risk nodes.
@@ 1 Introduction
The United States is currently facing an alarming escalation in aviation accidents. In the first six weeks of 2025 alone, there have been over 30 commercial aviation incidents/accidents, including four major plane crashes that have tragically claimed 85 lives (Figure 1). This surge in accidents is unprecedented, especially considering that prior to 2025, the most recent fatal crash involving a U.S. airliner dated back in 2009 [1]. This disturbing trend has severely undermined public confidence in air travel. This erosion of trust not only deters passengers but also poses significant economic threats to the aviation industry, which is already grappling with financial instability. Immediate and robust research and development efforts are imperative to enhance safety protocols, rectify systemic deficiencies, and restore public trust in air travel. The safety of airport surface operations remains a critical challenge, particularly with the increasing complexity and traffic volume of modern airports [2]. The safety of airport surface operations encompasses both runway and taxiway environments, where the movement of aircraft and ground vehicles occurs in close proximity [3, 4]. The complexity of surface operations stems from the diverse and dynamic nature of airport environments. The Flight Safety Foundation (FSF) reported that 30% of commercial aviation accidents between 1995 and 2008 were runway-related, resulting in 973 fatalities [5].
@@
Runway incursions, defined as the unauthorized presence of an aircraft, vehicle, or person on a runway [7], represent one of the most significant threats to safe airport operations, where a substantial portion attributed to human error in communication between pilots and controllers. Runway incursions can be attributed to pilot deviations, operational incidents, and vehicle-related anomalies, collectively challenging the effectiveness of conventional safety protocols [8]. Federal Aviation Administration (FAA) recorded 7,864 runway incursions from 2012 and July 2017, with 63% (4,947 incidents) attributed to pilot deviations, 20% (1,529 incidents) to air traffic control (ATC) errors, and 17% (1,388 incidents) to vehicle or pedestrian deviations [9]. In 2024, FAA reported a total of 1,757 runway incursions, indicating a persistent safety concern [10]. Additionally, FAA identified an annual occurrence rate of 0.200 Category A and B incursions per million operations in 2013, signaling the persistent risk posed by ground-based collisions [11]. One recent runway incursion case is the Henada airport runway collision happened on January 2nd 2024, where an Airbus A350 collided with a Japan Coast Guard airplane during landing. The collision resulted in both aircraft catching fire. Although all 379 occupants of the Japan Airlines flight were safely evacuated, five of the six crew members on the Coast Guard aircraft perished. The accident was due to the human error, citing miscommunication and misunderstanding between the Coast Guard pilot and air traffic control as a primary factor [12].
@@
While runway incursions have historically garnered significant attention due to their potential for catastrophic outcomes, the number of taxiway collisions is also increasing and gaining public attention. Taxiway collisions occur when aircraft or vehicles inadvertently occupy the same space on the ground, lead to significant safety hazards, operational disruptions, and financial losses. Studies have shown that areas with higher taxiway occupancy rates are more prone to collisions, especially during peak operational periods. For instance, research indicates that increased workload during busy times can elevate collision risks, making it imperative to focus on these high-density areas, or potential collision spots, for effective risk management [13]. To address taxiway collision, the Surface Safety Risk Index (SSRI) is developed by FAA, which assigns risk weights to various outcomes, such as aircraft damage or injuries, based on their proximity to potential fatalities. Another recent taxiway collision case happened on September 10th 2024, at the Hartsfield-Jackson Atlanta International Airport (KATL) involved an Airbus A350’s wing striking the tail of a CRJ-900 regional jet during taxiing, resulting in significant damage to both aircraft [14].
@@
These statistics and accident cases underscore the persistent challenges associated with both runway incursions and taxiway collisions, highlighting the critical need for enhancing surface monitoring and risk assessment capabilities. Efforts have been made towards this direction. A comprehensive, technology-driven efforts have been made towards enhancing runway safety, such as integrating advanced surveillance, alerting systems, and infrastructural improvements. Runway Status Lights (RWSL) serve as a critical safety layer by utilizing real-time surveillance data to alert aircrews and vehicle operators to hazardous runway conditions, while surface surveillance systems such as ASDE‑X or ASSC provide air traffic controllers with highly accurate, integrated displays of surface movements. Enhancements like Taxiway Arrival Prediction (ATAP) further refine these systems by alerting controllers when aircraft deviate from their assigned paths, thereby reducing misalignment risks [15]. However, neither ASDE-X nor ASSC considers the input from the ATC radio communication signals, making it hard to check for pilot-ATC communication induced risks. The rapid advancement of artificial intelligence (AI) paves the way to understand and extract key information from the voice communications through machine intelligence. Natural Language Processing (NLP), as a sub-field of AI, focuses on understanding and analyzing the contextual meaning of human spoken language, is thus considered in this research work. Specifically, the NLP model working on understanding the voice communication transcripts is also known as Automatic Speech Recognition (ASR). On the tactical level, extracting keywords from the textual format transcripts is one of the major interests. Information retrieval takes keywords and useful contexts from ASR converted text using either token-based or contextual-based language model. This process is known as Named Entity Recognition (NER) where entities are the extracted keywords. A detailed review of the various usage of NER is given in Section 2.1.
@@
Despite the demonstrated effectiveness of NLP to other domains, air traffic control is heavily regulated by communication rules and standards to ensure clarity, consistency, and safety in all ATC interactions. For instance, the air traffic control manual FAA JO 7110.65 defines the standardized phraseology and terminology that controllers are required to use during pilot-ATC communications [16]. [17] provides the word and phrase contractions used by all of the air traffic management related parties. This leads to the question: how can these heavily regulated ATC rules be incorporated into a data-driven NLP framework? On the other side, although English is the standard language used for air traffic control, multilingualism in aviation communications warrants careful consideration because the inherently international nature of air travel means that pilots and controllers often operate with diverse linguistic backgrounds, miscommunications stemming from accents, code-switching, and non-standard phraseology can lead to critical errors that compromise safety [18, 19]. This posts the challenge: how can the model be ensured to be multilingual for potential usage?. Moreover, recent advancement of large language models (LLMs) in NLP enables the capability of performing various tasks through prompt engineering. These models can also be adapted to different domains using either fine-tuning or domain-specific Retrieval-Augmented Generation (RAG) [20]. However, these LLM-based solutions require a significant amount of computer power during both the training and inference stages [20, 21, 22], which are extremely critical for onboard applications in near real-time. In [20], the llama2-based tuning process requires hundreds of gigabyte (GB) of CPU and GPU memory to fine-tune, and requires at least 25 GB of GPU memory during machine learning inference. Impressively, [21] mentioned the time required for pre-training of SafeAeroBERT takes over two months. This leads to the third focus: how can we ensure the model is of suitable size for efficient real-time information extraction on-prem? Lastly, the risk modeling of surface collision probability presents the challenge: how can the problem be formulated in a spatiotemporal sense, based on the estimated travel time distribution to the potential collision spots, which can utilize the output from the novel rule-enhanced machine learning model?
@@
We created our own ATC communication transcript dataset and annotated into NER training format, as well as our approach of encoding ATC rules as regular expressions.
@@
We provide the approach of incorporating the ATC-rules into NER training, evaluate and compare different setups and token-level embedding models and show the time and space complexity between setups.
@@
We propose the contextual similarity based search to link the destination ATC phraseology to the node names in NASA FACET layout maps.
@@
We propose the link travel time based airport surface collision risk assessment formulation, and show effectiveness on two recently occurred real-world accident scenarios.
The rest of the paper is organized as follows: Section 2 gives a literature review of related literature, where a review of the usage of language AI in aviation Section 2.1, along with a review of surface collision risk model is provided in Section 2.2. Section 3 discusses the proposed methodology, which is composed of two sub-modules, the ATC rule-enhanced learning module Section 3.1 and the risk formulation in Section 3.2. We demonstrate our approach with two case studies in Section 4, both of which occurred recently. Section 5 concludes the paper.
@@ 2.1 Language AI in Aviation
Natural language processing (NLP) applies computational techniques to learn, understand, and generate human language, evolving from early rule‐based and symbolic approaches to modern statistical and deep learning methods [24, 25]. Early systems relied on handcrafted grammars and rules, but the availability of large annotated corpora and advances in machine learning spurred breakthroughs in tasks such as machine translation [26, 27], where phrase‐based models have gradually been supplanted by neural network architectures [28, 29], as well as in speech recognition, dialogue systems, and sentiment analysis [30, 31]. These advancements have enabled real-world applications like conversational agents and real-time translation services, yet challenges remain in modeling complex semantic nuances and extending robust NLP capabilities to low-resource languages [32]. Two major research directions in NLP for Aviation safety, (a) aviation safety analysis on incident/accident reports; (b) air traffic control (ATC) communications transcripts analysis [33]. These research directions are essential for enhancing safety, operational efficiency, and decision-making processes. By leveraging NLP, we can achieve more accurate and timely analysis of safety data, improve the clarity and effectiveness of communication between pilots and air traffic controllers, and better manage the growing complexity of air traffic systems [34, 35].
@@
The key research direction of NLP usage in aviation safety reports include root cause analysis and critical factors identification related to aviation incidents/accidents, given risk assessment insights to inform risk management strategies and proactive operations. Topic modeling and pattern analysis is also a way to uncover and understand the latent behavior and common pattern of multiple incidents/accidents, such that operators can identify emerging risks with similar patterns from historical reports. U.S. National Aeronautics and Space Administration (NASA) Aviation Safety Report System (ASRS) database and the U.S. National Transportation Safety Board (NTSB) are two commonly used database for aviation safety/accident reports. Machine learning classification models such as SVM, CNN, RNN, LSTM are adopted in the literature. For instance, [36] proposes a method for incident root course analysis using the weakly supervised semantic lexicon learning and support vector machine (SVM) for root cause identification from ASRS. On the technical level, researchers propose various methods to identify risk-related causes from flight status [37, 38, 39], human factors [40, 41, 42], spatial-temporal relations [43], and anomalies [44, 45]. However, contextual ambiguity and multilingual scenarios usually leads to misinterpretations or errors in these studies.
@@
For communication transcripts, Automatic Speech Recognition (ASR), information retrieval, error detection, and speaker classification are major research directions [33]. ASR processes audio signals to identify words and phrases spoken by a human speaker, transcribing them into a textual format that can be analyzed and used by various applications [34, 46]. Information retrieval takes keywords and useful contexts from ASR converted text using either token-based or contextual-based language model. This process is known as Named Entity Recognition (NER) where entities are the extracted keywords. In air traffic communication transcripts understanding, the ability to accurately extract ATC domain‐specific entities (such as aircraft identifiers, altitudes, call signs, and route information) is critical for both operational safety and real-time decision making, where flight callsigns, and the intended destination are of critical interests [47, 36, 48]. NER-based information extraction communication transcripts is the key step of building a ATC deviation warning system, where the computer can determine of the aircraft is showing the expected behavior based on the extracted keywords and real-time location.
@@
NER methodologies in aviation communication transcript understanding have evolved from rule-based systems to deep learning-based models. Early work in aviation NER relied on expert-crafted rules using regular expressions and domain-specific lexicons to capture the complex, standardized language of ATC communications. These systems offer high interpretability and precision under controlled conditions; however, their rigidity limits adaptability to variations in speaker accents, noise levels, and spontaneous deviations. Statistical techniques such as Conditional Random Fields (CRFs) and Support Vector Machines (SVMs) are used to overcome the inflexibility of rule-based methods. These models effectively capture contextual dependencies but demand significant feature engineering (i.e., parts-of-speech, orthographic and positional features). Recent advances have seen a marked shift towards deep learning techniques for NER in aviation transcripts. These models automatically learn contextual representations from raw input text, reducing the need for extensive manual feature engineering [60]. [34] employed Mozilla’s Deep Speech implementation in combination with the Spacy library’s NER module to process ATC communication transcripts. Their system leverages deep learning to capture both the sequential context of spoken instructions and the unique syntactic patterns of ATC language. Moreover, transformer-based models are emerging as strong candidates due to their capacity for modeling long-range dependencies and handling variable-length sequences, which is particularly beneficial given the noisy and often rapidly spoken nature of ATC communications. Pre-trained language models (such as BERT or its derivatives) can be fine-tuned on domain-specific datasets, allowing these models to adapt to the peculiarities of aviation language switching phenomena [61]. The recent advancement of large language models (LLMs) has enabled the possibility of unified solutions of the above two directions. LLMs can achieve multiple tasks from document understanding to communication transcripts extraction thanks to prompt engineering, and are able to generated to different domains using either fine-tuning or Retrieval-Augmented Generation (RAG) [20]. However, these LLM-based solutions require significant amount of computer power during both the training and inference stage [20, 21, 22], which are drawbacks for onboard applications in near real-time. In [20], the Llama2-based tuning process requires hundreds of gigabyte (GB) of CPU and GPU memory to fine-tune, and requires at least 25 GB of GPU memory for machine learning inference. Impressively, [21] mentioned the time required for pre-training of SafeAeroBERT takes over two months. The computational demands associated with the use of generative AI render these models impractical for real-time aviation decision support in the current stage.
@@
As a summary, these approach offer distinct advantages, (a) rule-based methods provide high precision in controlled settings; (b) statistical machine learning methods offer robust performance with careful feature engineering; (c) deep learning techniques deliver state-of-the-art results in complex, real-world conditions; (d) LLMs are powerful and adaptable but requires substantial computational effort. The research trends indicate a clear movement toward end-to-end, hybrid systems that not only address the inherent challenges of noisy, rapid, and accented ATC communications but also integrate domain-specific knowledge to enhance safety and operational efficiency. These systems often incorporate expert-defined rules to pre-filter the input or to post-process the output of statistical or deep learning models. This approach not only leverages the precision of expert rules but also benefits from the adaptability and generalization capabilities of modern neural architectures.
@@
In this work, we address the aforementioned issues by first build our own data processing pipeline to process open source communication audios and scanned documents using the state-of-the-art speech-to-text engine. Then, we propose a hybrid ATC domain specific rule-enhanced NER to extract key information (i.e., callsign and destination intent) from the pilot-ATC communication transcripts, along with a post-prediction heuristic rule override process to further boost the recognition performance. The performance improvement of such post-recognition processing technique on intent learning is also highlighted in [47]. We examine the usage of various token-level embedding models with specifications such as multilingual support and generative power. Finally, we match the extracted entities with the node names from the NASA FACET node-link graph, based on embedding similarities.

### v2

@@ Abstract
Surface movement collision risk is critical for airport safety. These models play a vital role in identifying and mitigating potential hazards during airport ground operations by providing warnings of near-miss incidents, thereby reducing the risk of accidents that could jeopardize human lives and financial assets. However, existing models, developed decades ago, have not fully integrated recent advancements in machine intelligence, where incorporating additional functionalities presents promising opportunities for improved risk assessment. This work provides a feasible solution to the existing airport surface safety monitoring capabilities (i.e., Airport Surface Surveillance Capability (ASSC)), namely language AI-based voice communication understanding for collision risk assessment. The proposed framework consists of two major parts, (a) rule-enhanced Named Entity Recognition (NER); (b) surface collision risk modeling. NER module generates information tables by processing voice communication transcripts, which serve as references for producing potential taxi plans and calculating the surface movement collision risk. We first collect and annotate our dataset based on open-sourced video recordings and safety investigation reports. Additionally, we refer to FAA Order JO 7110.65W and FAA Order JO 7340.2N to get the list of heuristic rules and phase contractions of communication between the pilot and the Air Traffic Controller (ATCo). Then, we propose the novel ATC Rule-Enhanced NER method, which integrates the heuristic rules into the model training and inference stages, resulting in a hybrid rule-based NER model. We show the effectiveness of this hybrid approach by comparing different setups with different token-level embedding models. For the risk modeling, we adopt the node-link airport layout graph from NASA FACET and model the aircraft taxi speed at each link as a log-normal distribution and derive the total taxi time distribution. Then, we propose a spatiotemporal formulation of the risk probability of two aircraft moving across potential collision nodes during ground movement. Furthermore, we propose the real-time implementation of such a method to obtain the lead time, with a comparison with a Petri-Net based method. We show the effectiveness of our approach through case studies, (a) the Haneda airport runway collision accident happened in January 2024; (b) the KATL taxiway collision happened in September 2024; (c) the Tenerife airport disaster in March 1977. We show that, by understanding the pilot-ATC communication transcripts and analyzing surface movement patterns, the proposed model estimates the surface movement collision probability within machine processing time, thus enabling proactive measures to possible collisions at a certain node, which improves airport safety. A study on validating the log-normal assumption of aircraft taxi speed distributions is also given. We provide the link to code and data repository HERE.
@@
We introduce a novel framework to integrate natural language processing with surface collision risk modeling to enhance aviation safety.
@@
Surface movements are modeled with log-normal distributions and employ node-link graph structures to estimate spatiotemporal collision probabilities.
Real-time risk assessment at overlapping nodes with different warning thresholds and lead time analysis.
We validate the log-normal link travel speed assumptions by conducting data analysis and statistical tests on ASDE-X ground movement data.
Three case studies, the Haneda runway collision, the KATL taxiway collision, and the Tenerife airport disaster, demonstrate the effectiveness in detecting high-risk nodes.
@@ 1 Introduction
The United States is currently facing an alarming escalation in aviation accidents. In the first six weeks of 2025 alone, there have been over 30 commercial aviation incidents/accidents, including four major plane crashes that have tragically claimed 85 lives (Figure 1). This surge in accidents is unprecedented, especially considering that prior to 2025, the most recent fatal crash involving a U.S. airliner dated back in 2009 [1]. This disturbing trend not only deters passengers but also poses significant economic threats to the industry, which is already grappling with financial instability. Immediate and robust research and development efforts are imperative to enhance safety protocols, rectify systemic deficiencies, and restore public trust in air travel. The safety of airport surface operations remains a critical challenge, particularly with the increasing complexity and traffic volume of modern airports [2]. The safety of airport surface operations encompasses both runway and taxiway environments, where the movement of aircraft and ground vehicles occurs in close proximity [3, 4]. The complexity of surface operations stems from the diverse and dynamic nature of airport environments. For instance, the Flight Safety Foundation (FSF) reported that 30% of commercial aviation accidents between 1995 and 2008 were runway-related, resulting in 973 fatalities [5].
@@
Runway incursions, defined as the unauthorized presence of an aircraft, vehicle, or person on a runway [7], represent one of the most significant threats to safe airport operations, where a substantial portion is attributed to human error in communication between pilots and controllers. Runway incursions can be attributed to pilot deviations, operational incidents, and vehicle-related anomalies, collectively challenging the effectiveness of conventional safety protocols [8]. Federal Aviation Administration (FAA) recorded 7,864 runway incursions from 2012 and July 2017, with 63% (4,947 incidents) attributed to pilot deviations, 20% (1,529 incidents) to air traffic control (ATC) errors, and 17% (1,388 incidents) to vehicle or pedestrian deviations [9]. In 2024, FAA reported a total of 1,757 runway incursions, indicating a persistent safety concern [10]. Additionally, FAA identified an annual occurrence rate of 0.200 Category A and B incursions per million operations in 2013, signaling the persistent risk posed by ground-based collisions [11]. One recent runway incursion case is the Haneda airport runway collision happened on January 2nd 2024, where an Airbus A350 collided with a Japan Coast Guard airplane during landing. The collision resulted in both aircraft catching fire. Although all 379 occupants of the Japan Airlines flight were safely evacuated, five of the six crew members on the Coast Guard aircraft perished. The accident was due to human error, citing miscommunication and misunderstanding between the Coast Guard pilot and air traffic control as a primary factor [12].
@@
While runway incursions have historically garnered significant attention due to their potential for catastrophic outcomes, the number of taxiway collisions is also increasing and gaining public attention. Taxiway collisions occur when aircraft or vehicles inadvertently occupy the same space on the ground, leading to significant safety hazards, operational disruptions, and financial losses. Studies have shown that areas with higher taxiway occupancy rates are more prone to collisions, especially during peak operational periods. For instance, research indicates that increased workload during busy times can elevate collision risks, making it imperative to focus on these high-density areas, or potential collision spots, for effective risk management [13]. To address taxiway collisions, the Surface Safety Risk Index (SSRI) is developed by the FAA, which assigns risk weights to various outcomes, such as aircraft damage or injuries, based on their proximity to potential fatalities. Another recent taxiway collision case happened on September 10th 2024, at the Hartsfield-Jackson Atlanta International Airport (KATL) involved an Airbus A350’s wing striking the tail of a CRJ-900 regional jet during taxiing, resulting in significant damage to both aircraft [14].
@@
These statistics and accident cases underscore the persistent challenges associated with both runway incursions and taxiway collisions, highlighting the critical need for enhancing surface monitoring and risk assessment capabilities. Effort has been made towards this direction. A comprehensive, technology-driven efforts have been made towards enhancing runway safety, such as integrating advanced surveillance, alerting systems, and infrastructural improvements. Runway Status Lights (RWSL) serve as a critical safety layer by utilizing real-time surveillance data to alert aircrews and vehicle operators to hazardous runway conditions, while surface surveillance systems such as ASDE‑X or ASSC provide air traffic controllers with highly accurate, integrated displays of surface movements. Enhancements like Taxiway Arrival Prediction (TAP) further refine these systems by alerting controllers when aircraft deviate from their assigned paths, thereby reducing misalignment risks [15]. However, neither ASDE-X nor ASSC considers the input from the ATC radio communication signals, making it hard to check for pilot-ATC communication induced risks. The rapid advancement of artificial intelligence (AI) paves the way to understand and extract key information from the voice communications through machine intelligence. Natural Language Processing (NLP), as a sub-field of AI, focuses on understanding and analyzing the contextual meaning of human spoken language, and is thus considered in this research work. Specifically, the NLP model working on understanding the voice communication transcripts is also known as Automatic Speech Recognition (ASR). On the tactical level, extracting keywords from the textual format transcripts is one of the major interests. Information retrieval takes keywords and useful contexts from ASR-converted text using either token-based or contextual-based language model. This process is known as Named Entity Recognition (NER) where entities are the extracted keywords. A detailed review of the various usage of NER is given in Section 2.1.
@@
Despite the demonstrated effectiveness of NLP to other domains, air traffic control is heavily regulated by communication rules and standards to ensure clarity, consistency, and safety in all ATC interactions. For instance, the air traffic control manual FAA JO 7110.65 defines the standardized phraseology and terminology that controllers are required to use during pilot-ATC communications [16]. [17] provides the word and phrase contractions used by all of the air traffic management related parties. This leads to the question: how can these heavily regulated ATC rules be incorporated into a data-driven NLP framework? On the other side, although English is the standard language used for air traffic control, multilingualism in aviation communications warrants careful consideration because the inherently international nature of air travel means that pilots and controllers often operate with diverse linguistic backgrounds, miscommunications stemming from accents, code-switching, and non-standard phraseology can lead to critical errors that compromise safety [18, 19]. This poses the challenge: how can the model be ensured to be multilingual for potential usage?. Moreover, recent advancements of large language models (LLMs) in NLP enables the capability of performing various tasks through prompt engineering. These models can also be adapted to different domains using either fine-tuning or domain-specific Retrieval-Augmented Generation (RAG) [20]. However, these LLM-based solutions require a significant amount of computer power during both the training and inference stages [20, 21, 22], which are extremely critical for onboard applications in near real-time. In [20], the llama2-based fine-tuning process requires hundreds of gigabyte (GB) of CPU and GPU memory, and at least 25 GB of GPU memory during inference. Impressively, [21] mentioned the time required for pre-training of SafeAeroBERT takes over two months. This leads to the third focus: how can we ensure the model is of suitable size for efficient real-time information extraction on-prem? Lastly, the existing collision risk models fail to consider the spatial layout of the airport in a spatiotemporal fashion [13], which presents the challenge: how can the ground risk be formulated in a spatiotemporal sense, based on the estimated travel time distribution to the potential collision spots, which can utilize the output from the novel rule-enhanced machine learning model?
@@
We created our own ATC communication transcript dataset and annotated in NER training format, as well as our approach of encoding ATC rules as regular expressions.
@@
We provide the hybrid-learning approach of incorporating the ATC-rules into NER training. Through evaluating and comparing different setups and token-level embedding models, we show the complexity and sensitivity studies between setups.
@@
We propose the link travel speed based spatiotemporal airport ground collision risk formulation, and provide the real-time implementation of risk warning system.
@@
We show effectiveness of the proposed framework through the reconstruction of three real-world accident scenarios.
The rest of the paper is organized as follows: Section 2 gives a literature review of related literature, where a review of the usage of language AI in aviation Section 2.1, along with a review of surface collision risk model is provided in Section 2.2. Section 3 discusses the proposed methodology, which is composed of two sub-modules, the ATC rule-enhanced learning module Section 3.1 and the risk formulation in Section 3.2. We demonstrate our approach with three case studies in Section 4. Section 5 concludes the paper.
@@ 2.1 Language AI in Aviation
Natural language processing (NLP) applies computational techniques to learn, understand, and generate human language, evolving from early rule‐based and symbolic approaches to modern statistical and deep learning methods [24, 25]. Early systems relied on handcrafted grammars and rules, but the availability of large annotated corpora and advances in machine learning spurred breakthroughs in tasks such as machine translation [26, 27], where phrase‐based models have gradually been supplanted by neural network architectures [28, 29], as well as in speech recognition, dialogue systems, and sentiment analysis [30, 31]. These advancements have enabled real-world applications like conversational agents and real-time translation services, yet challenges remain in modeling complex semantic nuances and extending robust NLP capabilities to low-resource languages [32]. Two major research directions in NLP for Aviation safety are, (a) aviation safety analysis on incident/accident reports; (b) air traffic control (ATC) communications transcripts analysis [33]. These research directions are essential for enhancing safety, operational efficiency, and decision-making processes. By leveraging NLP, we can achieve more accurate and timely analysis of safety data, improve the clarity and effectiveness of communication between pilots and air traffic controllers, and better manage the growing complexity of air traffic systems [34, 35].
@@
The key research direction of NLP usage in aviation safety reports include root cause analysis and critical factors identification related to aviation incidents/accidents, giving risk assessment insights to inform risk management strategies and proactive operations. Topic modeling and pattern analysis is also a way to uncover and understand the latent behavior and common pattern of multiple incidents/accidents, such that operators can identify emerging risks with similar patterns from historical reports. U.S. National Aeronautics and Space Administration (NASA) Aviation Safety Report System (ASRS) database and the U.S. National Transportation Safety Board (NTSB) are two commonly used databases for aviation safety/accident reports. Machine learning classification models such as SVM, CNN, RNN, LSTM are adopted in the literature. For instance, [36] proposes a method for incident root course analysis using the weakly supervised semantic lexicon learning and support vector machine (SVM) for root cause identification from ASRS. On the technical level, researchers propose various methods to identify risk-related causes from flight status [37, 38, 39], human factors [40, 41, 42], spatial-temporal relations [43], and anomalies [44, 45]. However, contextual ambiguity and multilingual scenarios usually lead to misinterpretations or errors in these studies.
@@
For communication transcripts, Automatic Speech Recognition (ASR), information retrieval, error detection, and speaker classification are major research directions [33]. ASR processes audio signals to identify words and phrases spoken by a human speaker, transcribing them into a textual format that can be analyzed and used by various applications [34, 46]. Information retrieval takes keywords and useful contexts from ASR converted text using either token-based or contextual-based language model. This process is known as Named Entity Recognition (NER) where entities are the extracted keywords. In air traffic communication transcripts understanding, the ability to accurately extract ATC domain‐specific entities (such as aircraft identifiers, altitudes, call signs, and route information) is critical for both operational safety and real-time decision making, where flight callsigns, and the intended destination are of critical interests [47, 36, 48]. NER-based information extraction communication transcripts are the key step of building a ATC deviation warning system, where the computer can determine of the aircraft is showing the expected behavior based on the extracted keywords and real-time location. Recent work also emphasizes that AI-driven advisory systems in ATC must also address explainability and trust calibration, which reveals the need for explanations varies with operational goals and explanation design should begin with controllers’ reasoning needs rather than researcher assumptions [49].
@@
NER methodologies in aviation communication transcript understanding have evolved from rule-based systems to deep learning-based models. Early work in aviation NER relied on expert-crafted rules using regular expressions and domain-specific lexicons to capture the complex, standardized language of ATC communications. These systems offer high interpretability and precision under controlled conditions; however, their rigidity limits adaptability to variations in speaker accents, noise levels, and spontaneous deviations. Statistical techniques such as Conditional Random Fields (CRFs) and Support Vector Machines (SVMs) are used to overcome the inflexibility of rule-based methods. These models effectively capture contextual dependencies but demand significant feature engineering (i.e., parts-of-speech, orthographic and positional features). Recent advances have seen a marked shift towards deep learning techniques for NER in aviation transcripts. These models automatically learn contextual representations from raw input text, reducing the need for extensive manual feature engineering [61]. [34] employed Mozilla’s Deep Speech implementation in combination with the Spacy library’s NER module to process ATC communication transcripts. Their system leverages deep learning to capture both the sequential context of spoken instructions and the unique syntactic patterns of ATC language. Moreover, transformer-based models are emerging as strong candidates due to their capacity for modeling long-range dependencies and handling variable-length sequences, which is particularly beneficial given the noisy and often rapidly spoken nature of ATC communications. Pre-trained language models (such as BERT or its derivatives) can be fine-tuned on domain-specific datasets, allowing these models to adapt to the peculiarities of aviation language switching phenomena [62]. The recent advancement of large language models (LLMs) has enabled the possibility of unified solutions of the above two directions. LLMs can achieve multiple tasks from document understanding to communication transcripts extraction thanks to prompt engineering, and are able to be generated to different domains using either fine-tuning or Retrieval-Augmented Generation (RAG) [20]. However, these LLM-based solutions require a significant amount of computer power during both the training and inference stage [20, 21, 22], which are drawbacks for onboard applications in near real-time. In [20], the Llama2-based tuning process requires hundreds of gigabytes (GBs) of CPU and GPU memory to fine-tune, and requires at least 25 GB of GPU memory for machine learning inference. Impressively, [21] mentioned the time required for pre-training of SafeAeroBERT takes over two months. The computational demands associated with the use of generative AI render these models impractical for real-time aviation decision support in the current stage.
@@
As a summary, these approaches offer distinct advantages, (a) rule-based methods provide high precision in controlled settings; (b) statistical machine learning methods offer robust performance with careful feature engineering; (c) deep learning techniques deliver state-of-the-art results in complex, real-world conditions; (d) LLMs are powerful and adaptable but require substantial computational effort. The research trends indicate a clear movement toward end-to-end, hybrid systems that not only address the inherent challenges of noisy, rapid, and accented ATC communications but also integrate domain-specific knowledge to enhance safety and operational efficiency. These systems often incorporate expert-defined rules to pre-filter the input or to post-process the output of statistical or deep learning models. This approach not only leverages the precision of expert rules but also benefits from the adaptability and generalization capabilities of modern neural architectures.
@@
In this work, we address the aforementioned issues by first building our own data processing pipeline to process open source communication audios and scanned documents using the state-of-the-art speech-to-text engine. Then, we propose a hybrid ATC domain specific rule-enhanced NER to extract key information (i.e., callsign and destination intent) from the pilot-ATC communication transcripts, along with a post-prediction heuristic rule override process to further boost recognition performance. The performance improvement of such post-recognition processing technique on intent learning is also highlighted in [47]. We examine the usage of various token-level embedding models with specifications such as multilingual support and generative power. Finally, we match the extracted entities with the node names from the NASA FACET node-link graph, based on embedding similarities.